Chapter 7 Supplemental Practice Problems Mallu 6:7 2 number of standard deviations between x (variable of interest) and mu
Answer #1

α. Ρ(x < 113.25) = Φ(113.5 – 114)/7) = Φ(-0.07) = 1 - Φ(0.07) = 0.4721 6. Ρ(r> 113.25) = 1 – P(x < 113.25) = 1 – 0.4721 = 0.5

where (2) = cdf of standard normal distribution, 0(-)=1-02) and since z is continuous random variable so P(Z < 2) = P(Z < 2)